OpenStudy (anonymous):

PLEASE HELP will medal which of the following statements accurately compares f(x) and g(x)? choose all that apply. f(x)=30(1.2)^x A the growth factor of g(x) is less than the growth factor of f(x) B the growth factor of f(x)is less than the growth factor of g(x) C the initial values of f(x) and g(x) are the same D the decay factor of f(x) is less than the decay factor of g(x)

OpenStudy (anonymous):

table

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(x)=30(1.2)^x f(1)=30(1.2)^1 ... replace every x with 1 f(1) = ??

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

so is f(x) growing or decaying?

OpenStudy (anonymous):

growing?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

is it growing faster or slower than g(x)?

OpenStudy (anonymous):

slower

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so thats b right

OpenStudy (danjs):

i am late to the game

OpenStudy (danjs):

The site needs to upgrade servers badly.

OpenStudy (anonymous):

there should still be atleast one more answer

jimthompson5910 (jim_thompson5910):

yeah there are 2 answers

OpenStudy (anonymous):

is the other answer d

jimthompson5910 (jim_thompson5910):

do we have decay going on?

OpenStudy (anonymous):

yes?

jimthompson5910 (jim_thompson5910):

you sure?

jimthompson5910 (jim_thompson5910):

you said there was growth though

OpenStudy (anonymous):

oh ok then no

OpenStudy (anonymous):

is it c

jimthompson5910 (jim_thompson5910):

yeah notice how f(0) = 30 as well

OpenStudy (anonymous):

yeah so the final answers are b and c right

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok thank you